OpenStudy (anonymous):

okay so with exponential functions and looking at g(x)=2^x when its negative where does the 1/2 come from

OpenStudy (anonymous):

can you rephrase your question; I don't understand

OpenStudy (anonymous):

Do you mean the slope?

OpenStudy (anonymous):

g(x)=2^x the base of 2 is a fixed number and the x is your variable. Let's look more closely at the function g(x) = 2x. To evaluate this function, we operate as usual, picking values of x, plugging them in, and simplifying for the answers. But to evaluate 2x, we need to remember how exponents work. In particular, we need to remember that negative exponents mean "put the base on the other side of the fraction line". this is what the notes gave me

OpenStudy (anonymous):

i think i understand the question

OpenStudy (anonymous):

you mean when x is negative.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

So, a^-n = 1/a^n, for example.

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

so don't forget that \[b^{-n}=\frac{1}{b^n}\] so for example if you have \[2^x\] hten \[2^{-1}=\frac{1}{2}\] and \[2^{-2}=\frac{1}{2^2}=\frac{1}{4}\]

OpenStudy (anonymous):

thats right but that doesnt make sense on why the 1 is in the front and the variable is positive

OpenStudy (anonymous):

a negative number in the exponent does not make the number negative

OpenStudy (anonymous):

think of powers of ten

OpenStudy (anonymous):

so your saying that when the exponet is negative you have to take the recipocal of it?

OpenStudy (anonymous):

\[2^x = 1/(2^-n)\]

OpenStudy (anonymous):

\[10^3=1000\] \[10^2=100\] \[10^1=10\] \[10^0=1\] \[10^{-1}=0.1\] \[10^{-2}=0.01\] \[10^{-3}=0.001\] etc

OpenStudy (anonymous):

Wait, sorry! \[2^x = 1/(2^-x)\]

OpenStudy (anonymous):

yes the negative exponent means take the reciprocal

OpenStudy (anonymous):

@razzles 2^{-n}

OpenStudy (anonymous):

math is weird but not that weird. if you want the number to be negative, stick a minus sign in front

OpenStudy (anonymous):

\[2^{-2}=\frac{1}{4}\] \[-2^{-2}=-\frac{1}{4}\]

OpenStudy (anonymous):

Law of exponents: \[x ^{-a} = 1/ x ^{a}\]

OpenStudy (anonymous):

so with the negative numbers you move the decimal one place to the left?

OpenStudy (anonymous):

how are you getting 1/4 from 2^-2

OpenStudy (anonymous):

yes im this slow...

OpenStudy (anonymous):

got it now :D thanks guys